OpenStudy (davidusa):

CALCULUS HELP PLEASE. How do you determine whether this series (I'll write in question) diverges or converges?

OpenStudy (davidusa):

ill write it now

OpenStudy (anonymous):

alright bro

OpenStudy (davidusa):

summation of n from 1 to infinity of (arctan n)/(n^4)

OpenStudy (davidusa):

\[\sum_{n=1}^{\infty}\frac{ \tan^{-1} n }{ n^4 }\]

OpenStudy (davidusa):

There it is in real calculus form with all the signs and stuff.

OpenStudy (davidusa):

What test do you use? Ratio test? The Integral Test? The Direct Comparison Test? The Limit Comparison Test?

OpenStudy (davidusa):

And how do you do it? I don't need a complete explanation but at least explain it a little.

OpenStudy (anonymous):

bro it converges when you compare it with the p-series 1/n^4 since p = 4 it converges using i believe the limit comparison test..try that

OpenStudy (davidusa):

Well it diverges it says in the answer key... but i dont know how to get it. thanks for trying though.

OpenStudy (anonymous):

wait really? That's soo weird hmm..idk man i tried lol

OpenStudy (davidusa):

thanks for trying... i mean my sister says it diverges she's pretty sure. she's better than me at calculus, she's 3 years older than me. (she's in 10th grade im in 6th)

OpenStudy (anonymous):

wait what?! Bro ur in 6th grade?! how are you..omg i have full respect for u

OpenStudy (davidusa):

lol

OpenStudy (davidusa):

im 11 it says on my profile page. im 11 6th grade

OpenStudy (davidusa):

yea i know im a little young

OpenStudy (davidusa):

uhhh.... DUDE? u still there?